### v1

# CANDLE: Iterative Conceptualization and Instantiation Distillation from Large Language Models for Commonsense Reasoning

Abstract
The sequential process of conceptualization and instantiation is essential to generalizable commonsense reasoning as it allows the application of existing knowledge to unfamiliar scenarios. However, existing works tend to undervalue the step of instantiation and heavily rely on pre-built concept taxonomies and human annotations to collect both types of knowledge, resulting in a lack of instantiated knowledge to complete reasoning, high cost, and limited scalability. To tackle these challenges, we introduce CANDLE, a distillation framework that iteratively performs contextualized conceptualization and instantiation over commonsense knowledge bases by instructing large language models to generate both types of knowledge with critic filtering. By applying CANDLE to ATOMIC, we construct a comprehensive knowledge base comprising six million conceptualizations and instantiated commonsense knowledge triples. Both types of knowledge are firmly rooted in the original ATOMIC dataset, and intrinsic evaluations demonstrate their exceptional quality and diversity. Empirical results indicate that distilling CANDLE on student models provides benefits across four downstream tasks. Our code, data, and models are publicly available at https://github.com/HKUST-KnowComp/CANDLE.
1 Introduction
Commonsense reasoning refers to the cognitive ability to make logical inferences and draw conclusions based on general knowledge and understanding of the world that is typically shared among individuals Davis (2014); Mueller (2014). However, a longstanding challenge is its generalizability, as it often necessitates applying knowledge and reasoning to novel situations and contexts beyond simple pattern recognition or memorizing all special cases Mortimer (1995); Banaji and Crowder (1989). One promising approach to address this is the chain of conceptualization Murphy (2004) and instantiation Anderson et al. (1976), which, akin to the process of conceptual induction and deduction in human reasoning Tenenbaum et al. (2011), involves conceptualizing instances derived from known commonsense knowledge and subsequently instantiating these concepts in new situations to obtain the knowledge required for downstream reasoning. For example, in Figure 1, one can first conceptualize exercising in the gym as a healthy lifestyle, and further instantiate it to go on a balanced diet. This process allows the derivation of a novel event, PersonX goes on a balanced diet, which may entail new commonsense knowledge when placed back in the original context. By possessing substantial knowledge to initiate the process of conceptualization and instantiation, one can extrapolate limited commonsense knowledge to a wide array of diverse scenarios.
Yet, replicating this fundamental ability on machines remains challenging due to the absence of both types of knowledge in widely used CommonSense Knowledge Bases (CSKBs; Sap et al., 2019a; Speer et al., 2017; Fang et al., 2021b, a). Language models that are fine-tuned on these CSKBs are, therefore, unable to effectively utilize such a chain for reasoning during inference. To compensate, various methods have been proposed for entity-level Durme et al. (2009); Song et al. (2011, 2015); Gong et al. (2016); He et al. (2020); Peng et al. (2022a) and event-level Chen et al. (2020); He et al. (2022a); Wang et al. (2023b) conceptualizations by matching against concept taxonomies like Probase Wu et al. (2012) and WordNet Miller (1995). Meanwhile, Allaway et al. (2023) address instantiation through controllable generation. However, several limitations still persist.
Firstly, the two-step chain of conceptualization and instantiation can significantly augment a CSKB by introducing a diverse range of instances not originally included, thereby expanding the CSKB’s semantic coverage. However, most existing works underestimate the importance of the second step, instantiation, by focusing solely on conceptualization and using the resulting abstract knowledge directly. Other studies that concentrate on instantiations either overlook the conceptualization step entirely or only retrieve instances from the original CSKB, failing to introduce novel entities and events. Secondly, most conceptualization methods heavily depend on matching instances with concepts in concept taxonomies, such as Probase and WordNet, which have a limited scope and lack contextual information. Consequently, the derived conceptualizations are constrained in scale by these taxonomies and are formulated without considering proper contextualization, necessitating further verification in the original context. Lastly, the chain of conceptualization and instantiation can easily bring more than two orders of magnitude of data on top of the original CSKB. However, current acquisition and verification methods for both steps heavily rely on human annotation, which can be extremely costly as the scale of the CSKB increases.
To address these gaps, we introduce CANDLE, a ConceptuAlization and INstantiation Distillation framework from Large Language ModEls (LLMs) to aid commonsense reasoning. Specifically, CANDLE marks the first to complete the chain of conceptualization and instantiation by instructing powerful LLMs to sequentially generate both types of knowledge based on concrete commonsense triples while carefully considering the original context throughout the process. We further alleviate the human annotation cost by employing two critic filtering models to eliminate low-quality generations. The instantiated knowledge, representing concrete commonsense knowledge again, can be fed back into CANDLE as input, iteratively augmenting the original CSKB significantly.
By applying CANDLE to ATOMIC Sap et al. (2019a), we construct a large-scale knowledge base comprising 6.18 million conceptualizations and instantiations from two powerful LLMs, ChatGPT OpenAI (2022) and LLAMA2 Touvron et al. (2023). We demonstrate the intrinsic efficacy of CANDLE through automatic and human evaluations, showcasing the superior quality and diversity of knowledge generated by CANDLE (Section 5.1). We further show the extrinsic benefits of CANDLE by leveraging the generated knowledge as complementary training data to distill student models that yield improvements across four downstream tasks, including CSKB conceptualization, generative commonsense inference, and zero-shot commonsense question answering (Section 5.2).
2 Related Works
2.1 Conceptualization and Instantiation
Conceptualization aims to abstract a set of entities or events into a general concept, thereby forming abstract commonsense knowledge within its original context Murphy (2004). Subsequently, instantiation ground the derived concept into other instances and events to introduce new commonsense knowledge. Existing works primarily focused on entity-level conceptualization Durme et al. (2009); Song et al. (2011, 2015); Liu et al. (2022); Peng et al. (2022a), with He et al. (2022a) pioneering the construction of an event conceptualization benchmark. They combined syntactic parsing, semantically heuristic matching, and human annotation to extract concepts for social events from WordNet Miller (1995) synsets and Probase Wu et al. (2012). Building upon this work, Wang et al. (2023b) proposed a semi-supervised framework for conceptualizing CSKBs and demonstrated that abstract knowledge can enhance commonsense inference modeling and commonsense question answering Wang et al. (2023a). Wang et al. (2023d) also extended the study by constructing an abstraction benchmark based on eventualities from ASER Hongming Zhang et al. (2022). Regarding instantiation, Allaway et al. (2023) introduced a controllable generative framework to identify valid instantiations for abstract knowledge automatically. However, none of the existing studies have fully completed the chain of conceptualization and instantiation, with each focusing on only one aspect. Human annotation is also frequently applied for data collection and verification, which is both expensive and limited in scalability. Additionally, the downstream benefits of instantiated commonsense knowledge have not been thoroughly explored, leaving a significant gap in improving commonsense reasoning models.
2.2 Commonsense Knowledge Distillation
Recent breakthroughs in LLMs OpenAI (2022, 2023) have led to numerous efforts in distilling commonsense knowledge into datasets for training performant student models. West et al. (2022); Sclar et al. (2022); Bhagavatula et al. (2023); West et al. (2023) followed the pipeline of symbolic knowledge distillation, which uses human-crafted prompts to extract specific types of knowledge from LLMs for training downstream models. He et al. (2022b) proposed an alternative that involves transferring distilled knowledge from a ranker to a retriever, resulting in a more consistent and robust commonsense generator. Chae et al. (2023) and Kim et al. (2023) focused on distilling conversational responses from LLMs to enhance dialogue agents with commonsense knowledge and high-quality rationales. In this paper, we share similar aspirations and propose a chain of distillation framework that sequentially obtains abstract and instantiated knowledge from powerful LLMs. Empirical results show that our framework offers more substantial downstream benefits than previous symbolic knowledge distillation methods.
3 Definitions and Datasets
We follow the definitions proposed by He et al. (2022a) and Wang et al. (2023b) to formulate conceptualization and instantiation. Denote the triples in the original CSKB as , where , , and are the set of heads, relations, and tails in the original CSKB. The objective of conceptualization is to form a conceptualized head event, denoted as , from the original head . This is achieved by linking a component to a concept , forming by replacing with . Consequently, abstract knowledge is formed by combining the conceptualized head event with the original relation and tail, represented by . In the next step, the goal of instantiation is to associate the concept with a new instance . This process enables the formation of new commonsense knowledge in the format of , where is obtained by replacing with . In this paper, we use ATOMIC Sap et al. (2019a) as the original CSKB , which contains 310K triples after dropping those with wildcards and 18,839 unique head events. AbstractATOMIC He et al. (2022a) is used as the source of instances for every head event .
4 CANDLE
This section introduces our CANDLE framework, illustrated in Figure 2. Our framework can be outlined in three steps: (1) Instruct ChatGPT to generate contextualized conceptualizations based on the triples in the original CSKB. (2) Instruct LLAMA2 to instantiate the conceptualizations obtained in Step 1. (3) Apply critic-filtering to the generations in both steps and close the loop by reintroducing the instantiations back to the CSKB.
4.1 Contextualized Conceptualization
Previous methods for collecting conceptualizations rely on heuristically matching instances against concepts from WordNet and Probase. However, they suffer from limited concept coverage, resulting in a lack of knowledge diversity after instantiation, and require additional verification to ensure that concept fits into the original context . To address both issues, we propose to utilize ChatGPT as a loose teacher to collect conceptualizations in a one-step inference manner. Following Brown et al. (2020) and West et al. (2022), we use a few-shot prompt to instruct ChatGPT:
| <TASK-PROMPT> |
| <EX-INP><EX-OUT> |
| <EX-INP><EX-OUT> |
| <EX-INP> |
where <TASK-PROMPT> is a task instruction that explains how to conceptualize an event and <EX-INP><EX-OUT> are human authored examples of conceptualizations for events sampled from ATOMIC. For each example, are included in the input, and is the output. Finally, we provide the N input as <EX-INP> and ask ChatGPT to generate the corresponding conceptualization as <EX-OUT>. This ensures that ChatGPT not only learns the relationship between instances and their conceptualizations but also performs such abstraction in a contextualized manner, ensuring the plausibility of the generated conceptualization within the original context . Additionally, we include several hints in the prompt, such as numbering the examples, generating concepts instead of hypernyms, and keeping the generated responses concise. In this paper, we set N = 11 and obtain 20 conceptualizations for every event .
4.2 Contextualized Instantiation
After conceptualizing all events, we proceed to instantiate them by instructing an LLM once again. We use LLAMA2-13B as the loose teacher model for instantiation to reduce the significant cost associated with instantiating numerous conceptualizations. Our pilot study demonstrates that it can also generate high-quality instantiation knowledge. We employ a similar prompt as described in Section 4.1, with the modification of replacing <TASK-PROMPT> with the explanation of instantiating a conceptualized event and changing <EX-INP><EX-OUT> to human-authored examples of instantiations for abstract commonsense knowledge triples. are included in the input and is the expected output. By learning from these examples, LLAMA2 is expected to generate the corresponding instantiation (<EX-OUT>) based on the given abstract knowledge triple (<EX-INP>). We also set N = 11 and produce only one instantiation for each conceptualized event due to the significant amount of conceptualizations obtained in the previous step. Appendix A provides more details regarding the distillation process.
4.3 Iterating with Critic Filtering
Following West et al. (2022), we use critic filtering models to eliminate low-quality generations from LLMs. Specifically, we utilize a DeBERTa-v3-large conceptualization discriminator, provided by Wang et al. (2023b), and VERA-T5-xxl, provided by Liu et al. (2023), to evaluate the quality of the generated conceptualizations and instantiations, respectively. We set an empirical threshold value to serve as the cutoff point for discarding generations with scores below . In Section 5.1, we present evaluations conducted to determine the optimal value for . For all downstream applications, we set . Post-filtering, the instantiated triples can be reintroduced as the input for conceptualizations again as they continue to represent concrete commonsense knowledge. This iterative process of conceptualization and instantiation forms a loop, which enables continuously augmenting a CSKB. In this paper, we execute the loop only once, but multiple iterations hold the promise of significantly enhancing the CSKB’s knowledge coverage.
| Corpus | Conceptualization | Instantiation | ||
| Size (Unq.)/K | Accept | Size (Unq.)/K | Accept | |
| AbsATM | 503.5 (31.22) | - | None | - |
| EXEM | 0.650 (0.650) | - | 25.12 (25.12) | - |
| CANDLE | 6,181 (853.5) | 82.6% | 6,181 (676.7) | 77.9% |
| (critic) | 4,002 (498.4) | 88.1% | 4,176 (512.7) | 84.4% |
| (critic) | 3,272 (382.2) | 93.5% | 3,098 (455.9) | 89.1% |
| (critic) | 2,137 (219.4) | 97.2% | 2,208 (382.1) | 94.5% |
5 Evaluations and Analysis
In this section, we evaluate CANDLE from both intrinsic and extrinsic perspectives. Intrinsically, we demonstrate the high quality and diversity of conceptualizations and instantiations generated by CANDLE (Section 5.1). Extrinsically, we explore the benefits by applying the distilled knowledge to downstream tasks (Section 5.2).
5.1 Distillation Evaluations
Statistics and Quality.
We present the statistics of CANDLE distillation on ATOMIC in Table 1, demonstrating that CANDLE surpasses other existing benchmarks in terms of scale and concept coverage. Even under a rigorous critic filtering threshold (), CANDLE maintains its leading position with the highest count of total and unique knowledge for both types. To assess the quality of the distilled knowledge, we recruit four expert annotators, who are graduate students with expertise in commonsense, to conduct human evaluations on the plausibility of the generated conceptualizations and instantiations. Specifically, we randomly select 3,000 abstract commonsense triples and 3,000 instantiated triples from the distilled knowledge set. The annotators are then asked to determine (1) the correctness of the conceptualizations and instantiations and (2) the plausibility of their formed triples. Only triples whose both parts are deemed plausible by all annotators are considered accepted. We analyze the ratio of accepted triples among the samples with different levels of critic filtering and report the results in Table 1. Our findings demonstrate that LLMs exhibit impressive abilities in conceptualization and instantiation, with initial plausibility rates of 82.6% and 77.9% for both types of knowledge, respectively. With critic filtering, we observe an expected improvement in plausibility by maximally 14.6% and 16.6%, underscoring the effectiveness of our filtering measures in ensuring the high quality of the distilled knowledge. Details about annotations are in Appendix D.
Conceptualization Diversity.
The process of abstracting an event into highly diverse conceptualizations plays a crucial role in CANDLE. It is of significant importance because the greater the diversity of conceptualizations, the broader the knowledge coverage becomes upon instantiation. This, in turn, enhances the overall knowledge coverage within the distillation process. To examine the diversity of the top 10,000 popular distilled conceptualizations, we obtain their hypernyms by matching against Probase Wu et al. (2012) and present a visualization in Figure 3. It reveals that our distilled conceptualizations demonstrate a high level of diversity across various categories, forming a comprehensive and intricate knowledge base.
Additionally, following Wang et al. (2023d), we calculate the percentage of unique abstract concepts using BLEU soft uniqueness Zhu et al. (2018); West et al. (2022). We define a concept, denoted as , as unique if , where represents all concepts that share the same head event and identified instance with in AbstractATOMIC. Here, 0.5 serves as an empirical threshold. Our distillation process yields 92.3% unique conceptualizations, indicating a significantly higher diversity than previous datasets.
| Model Type | Backbone Model / Method | Event Conceptualization | Triple Conceptualization | ||
| Validation | Testing | Validation | Testing | ||
| Pre-trained Langauge Models | RoBERTa-large 340M | 77.28 | 77.99 | 81.77 | 82.69 |
| DeBERTa-v3-large 435M | 78.02 | 78.27 | 82.18 | 82.96 | |
| GPT2-XL 1.5B | 53.71 | 56.10 | 47.65 | 47.21 | |
| PseudoReasoner (RoBERTa-large) | 78.33 | 78.91 | 79.69 | 80.27 | |
| PseudoReasoner (DeBERTa-v3-large) | 79.03 | 79.21 | 79.89 | 80.07 | |
| CAT (RoBERTa-large) 340M | 78.51 | 78.53 | 82.27 | 83.02 | |
| CAT (DeBERTa-v3-large) 435M | 79.55 | 79.39 | 82.88 | 83.52 | |
| Large Language Models | ChatGPT (openai/gpt-3.5-turbo) | 69.29 | 68.65 | 68.54 | 68.12 |
| + Five-shot Exemplars | 69.42 | 70.40 | 70.27 | 72.08 | |
| + Chain-of-thought | 74.82 | 72.32 | 71.48 | 72.85 | |
| LLAMA2 7B | 46.29 | 43.90 | 40.81 | 41.25 | |
| + Five-shot Exemplars | 47.92 | 44.89 | 74.67 | 76.80 | |
| LLAMA2 13B | 48.17 | 48.59 | 48.31 | 48.55 | |
| + Five-shot Exemplars | 49.29 | 49.90 | 80.67 | 82.08 | |
| Mistral-v0.1 7B | 46.29 | 43.90 | 58.09 | 58.07 | |
| + Five-shot Exemplars | 51.00 | 50.06 | 65.09 | 69.80 | |
| LLAMA2 (LoRA Fine-tuned) 7B | 75.80 | 76.27 | 79.89 | 82.15 | |
| Mistral-v0.1 (LoRA Fine-tuned) 7B | 75.71 | 76.76 | 79.59 | 80.35 | |
| VERA-T5 5B | 70.76 | 70.29 | 72.60 | 76.85 | |
| VERA-T5 (Fine-tuned) 5B | 75.69 | 76.21 | 80.13 | 81.25 | |
| CANDLE Distilled (Ours) | RoBERTa-large 340M | 80.69 | 80.99 | 83.11 | 84.50 |
| DeBERTa-v3-large 435M | 80.97 | 81.14 | 83.64 | 84.64 | |
| LLAMA2 (LoRA Fine-tuned) 7B | 77.48 | 78.27 | 81.68 | 83.40 | |
| Mistral-v0.1 (LoRA Fine-tuned) 7B | 77.77 | 78.29 | 81.95 | 82.54 | |
| VERA-T5 (Fine-tuned) 5B | 77.54 | 78.03 | 82.79 | 83.61 | |
5.2 Downstream Applications
In this section, we explore the downstream applications of CANDLE. By applying CANDLE to ATOMIC, the distilled conceptualizations and instantiations form a large-scale expansion of the original CSKB, which contains high-quality abstract and concrete commonsense knowledge. Leveraging both types of knowledge as supplementary training data, we enhance various commonsense reasoning models for downstream tasks. Specifically, we utilize distilled conceptualizations in two CSKB conceptualization tasks Wang et al. (2023b), while instantiations are used in generative commonsense inference (COMET; Bosselut et al., 2019) and zero-shot commonsense QA tasks Ma et al. (2021). Due to space constraints, for all tasks, please refer to: Appendix B for dataset descriptions and Appendix C for implementation details.
5.2.1 CSKB Conceptualization
Task Setup.
The CSKB conceptualization task evaluates a model’s ability to conceptualize a CSKB through two binary classification subtasks, which are crucial for performing CSKB conceptualization inference upon concept taxonomies He et al. (2022a); Wang et al. (2023b). The first subtask, event conceptualization, aims to determine whether can be correctly conceptualized using , where is derived by replacing an instance with its linked concept . The second subtask, triple conceptualization, aims to assess the plausibility of a conceptualized triple that represents abstract commonsense knowledge. Accuracy is used as the evaluation metric. Following Wang et al. (2023b), we use the AbstractATOMIC dataset provided by He et al. (2022a) as the evaluation benchmark, which extends ATOMIC Sap et al. (2019a) by including human-annotated conceptualizations and abstract commonsense triples. The annotated data is partitioned by following ATOMIC’s original split of head events.
To obtain our distilled models for these tasks, we first synthesize negative samples from CANDLE distilled conceptualizations. For event conceptualizations, a random concept from another head event without common words is selected as the negative candidate, while for triple conceptualization, a tail of another head event without common words under the same relation is selected. We then fine-tune language models on a balanced mixture of CANDLE distillation and synthesized negative samples to train two models, each serving as a pre-trained general discriminator in their respective task domain. These two models are subsequently fine-tuned on the training sets of AbstractATOMIC to fit into the benchmark, and their performances on the validation and test sets are reported.
| Training Data | Bleu-1 | Bleu-2 | Bleu-3 | Bleu-4 | METEOR | ROUGE-L | CIDEr | BERTScore | Human |
| Backbone: GPT2-XL Radford et al. (2019) 1.5B | |||||||||
| Zero-shot | 4.350 | 1.598 | 0.732 | 0.293 | 5.702 | 5.030 | 0.792 | 37.11 | 14.50 |
| ATOMIC | 45.72 | 29.18 | 21.12 | 16.15 | 29.97 | 49.69 | 64.61 | 76.09 | 70.50 |
| ATOMIC | 42.15 | 25.77 | 17.82 | 13.14 | 29.82 | 47.61 | 63.70 | 70.39 | 76.50 |
| ATOMIC-10X | 45.38 | 29.20 | 21.09 | 16.15 | 30.09 | 49.86 | 65.02 | 75.89 | 77.50 |
| AbstractATOMIC | 45.30 | 29.08 | 21.00 | 16.06 | 29.98 | 48.61 | 63.98 | 75.56 | 71.50 |
| CANDLE Distilled | 50.71 | 33.85 | 25.55 | 20.43 | 32.45 | 51.91 | 69.68 | 76.86 | 78.50 |
| Backbone: ChatGPT OpenAI (2022) (openai/gpt-3.5-turbo) | |||||||||
| Zero-shot | 11.82 | 4.258 | 1.891 | 0.926 | 13.87 | 13.73 | 4.350 | 49.28 | 78.50 |
| Five-shot | 26.32 | 12.50 | 7.160 | 4.415 | 18.60 | 24.65 | 8.313 | 58.69 | 81.00 |
| Chain-of-thought | 9.906 | 3.568 | 1.556 | 0.736 | 11.85 | 11.02 | 2.905 | 46.17 | 64.00 |
| Backbone: LLAMA2 Touvron et al. (2023) 7B | |||||||||
| Zero-shot | 18.26 | 7.453 | 3.594 | 1.945 | 15.90 | 20.28 | 8.872 | 48.23 | 48.50 |
| Five-shot | 31.22 | 16.87 | 9.767 | 5.989 | 19.74 | 27.67 | 17.83 | 58.41 | 65.50 |
| ATOMIC | 42.04 | 23.01 | 14.10 | 9.125 | 27.80 | 42.90 | 53.17 | 71.52 | 68.50 |
| ATOMIC | 41.07 | 22.46 | 13.62 | 8.619 | 27.74 | 42.42 | 53.28 | 71.77 | 74.00 |
| ATOMIC-10X | 42.07 | 23.08 | 14.14 | 9.198 | 28.14 | 42.75 | 53.69 | 71.93 | 76.50 |
| AbstractATOMIC | 42.78 | 23.64 | 14.58 | 9.471 | 27.74 | 42.55 | 53.12 | 71.51 | 71.00 |
| CANDLE Distilled | 43.86 | 24.40 | 15.12 | 10.00 | 28.36 | 43.86 | 54.25 | 72.94 | 79.50 |
Baselines.
We compare our distilled models against the following baselines. First, we report the performance of language models when fine-tuned on the training set of AbstractATOMIC, including RoBERTa-Large Liu et al. (2019), DeBERTa-V3-Large He et al. (2023), GPT-2 Radford et al. (2019), LLAMA2 Touvron et al. (2023), Mistral Jiang et al. (2023), and VERA Liu et al. (2023). For fine-tuning LLMs, we adopt LoRA Hu et al. (2022) due to computational power limitations. Additionally, we incorporate semi-supervised learning methods that leverage both human-annotated and unlabeled data in AbstractATOMIC, including PsuedoReasoner Fang et al. (2022) and CAT Wang et al. (2023b). Finally, we include the performance of prompting LLMs, including LLAMA2, Mistral, and ChatGPT. In addition to direct zero-shot prompting, we also experiment with augmenting their performance with five-shot exemplars Wei et al. (2023) and chain-of-thought reasoning Wei et al. (2022).
Results and Analysis.
The results are reported in Table 2. Among all baselines, CAT trained with DeBERTa-v3-large demonstrates superior performance for both tasks compared to others. Within the baselines that use LLMs, LLAMA and Mistral yield the best results after fine-tuning. It’s worth noting that prompting ChatGPT with additional exemplars and chain-of-thought also produces commendable performances while prompting LLAMA and Mistral fail in most scenarios. However, pre-training language models on CANDLE distilled conceptualizations leads to consistently improved results for both tasks. This is evident when comparing the performance of each model with its respective best-performing baseline using the same backbone. For instance, fine-tuned Mistral shows a significant improvement of 1.54% and 2.19% on two tasks compared to fine-tuning on AbstractATOMIC, respectively. Remarkably, the distilled DeBERTa-v3-large surpasses all baselines and achieves state-of-the-art performances. This can be attributed to conceptualizations distilled from CANDLE, which provide the model with a broader understanding of various events and their conceptualizations, thus enhancing its discriminatory capabilities.
5.2.2 Generative Commonsense Inference
Task Setup.
The task of generative commonsense inference modeling (COMET; Bosselut et al., 2019) asks the model to generate commonsense tails based on given head and relation inputs. It serves as a fundamental component for numerous high-level applications that necessitate commonsense reasoning, such as zero-shot commonsense question answering with self-talk Shwartz et al. (2020) and dynamic graph construction Bosselut et al. (2021), narrative reasoning Peng et al. (2022b), and dialogue generation Tu et al. (2022). Improving COMET can potentially benefit other domains that require commonsense understanding. Following Hwang et al. (2021), we use the full test set of ATOMIC as our evaluation benchmark, which comprises 34,689 triples across 23 distinct commonsense relations. To automatically evaluate our model, we utilize a range of metrics, including BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Lavie and Agarwal (2007), CIDEr Vedantam et al. (2015), and BERTScore Zhang et al. (2020). We also recruit four expert annotators to conduct an expert evaluation of the models’ outputs. They are asked to annotate the plausibility of 200 randomly selected commonsense triple generations under each setting, and the resulting plausibility rates are reported.
Similar to training distilled models in previous tasks, we first pre-train GPT2 and LLAMA2-7B on critic-filtered CANDLE instantiations, where each triple is concatenated into a sentence via natural language templates. Subsequently, we fine-tune these models on the training split of ATOMIC to fit them into the benchmark. Finally, we report their performances on the test set.
Baselines.
For baselines, we use GPT2 and LLAMA2-7B as the backbone models. We train them separately on the training sets of ATOMIC, ATOMIC, ATOMIC10X West et al. (2022), and AbstractATOMIC. These models are then fine-tuned on the training split of ATOMIC and evaluated on the test set. We also include their zero-shot prompting performances. LLAMA2 is additionally evaluated for its prompting performance with five-shot exemplars. We also incorporate ChatGPT as another powerful commonsense generation baseline and provide its performance in zero-shot, five-shot, and chain-of-thought prompting scenarios.
Results and Analysis.
The results are presented in Table 3. Among the baselines, models that are pre-trained on ATOMIC-10X demonstrate the highest expert acceptance rate, surpassing even those pre-trained on AbstractATOMIC. One possible reason is that ATOMIC-10X encompasses a broader range of commonsense relations that are more consistent with the testing set of ATOMIC. However, models that are pre-trained on CANDLE instantiations achieve the highest scores compared to baselines with the same backbone model. For example, the CANDLE distilled LLAMA-7B model exhibits a 1.01% increase in BERTScore and a 3.00% increase in expert-plausibility compared to the best baseline. Even when compared to ChatGPT, CANDLE distilled models outperform it in all automatic metrics while maintaining a high plausibility rate of around 80%. We also observe that CANDLE distilled models consistently outperform those pre-trained on ATOMIC-10X and AbstractATOMIC across all metrics. This highlights the superior benefits of using CANDLE distilled instantiations for training COMET, compared to traditional symbolic knowledge distillation methods or augmenting the original CSKB with abstract commonsense knowledge.
| Model/Method | CSKB | a-NLI | CSQA | PIQA | SIQA | WG | Avg. |
| Pre-trained Language Models | |||||||
| RoBERTa-L Liu et al. (2019) | - | 65.5 | 45.0 | 67.6 | 47.3 | 57.5 | 56.6 |
| DeBERTa-v3-L He et al. (2023) | - | 59.9 | 25.4 | 44.8 | 47.8 | 50.3 | 45.6 |
| Self-talk Shwartz et al. (2020) | - | - | 32.4 | 70.2 | 46.2 | 54.7 | - |
| SMLM Banerjee and Baral (2020) | * | 65.3 | 38.8 | - | 48.5 | - | - |
| COMET-DynGen Bosselut et al. (2021) | ATOMIC | - | - | - | 50.1 | - | - |
| MICO Su et al. (2022) | ATOMIC | - | 44.2 | - | 56.0 | - | - |
| STL-Adapter Kim et al. (2022) | ATOMIC | 71.3 | 66.5 | 71.1 | 64.4 | 60.3 | 66.7 |
| DeBERTa-v3-L (MR) Ma et al. (2021) | ATM10X | 75.1 | 71.6 | 79.0 | 59.7 | 71.7 | 71.4 |
| DeBERTa-v3-L (MR) Ma et al. (2021) | ATOMIC | 76.0 | 67.0 | 78.0 | 62.1 | 76.0 | 71.8 |
| CAR-DeBERTa-v3-L Wang et al. (2023a) | ATOMIC | 78.9 | 67.2 | 78.6 | 63.8 | 78.1 | 73.3 |
| CAR-DeBERTa-v3-L Wang et al. (2023a) | AbsATM | 79.6 | 69.3 | 78.6 | 64.0 | 78.2 | 73.9 |
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2 | 69.9 | 80.3 | 65.9 | 78.3 | 74.9 |
| Large Language Models | |||||||
| GPT-3.5 (text-davinci-003) | - | 61.8 | 68.9 | 67.8 | 68.0 | 60.7 | 65.4 |
| ChatGPT (gpt-3.5-turbo) | - | 69.3 | 74.5 | 75.1 | 69.5 | 62.8 | 70.2 |
| + Chain-of-thought | - | 70.5 | 75.5 | 79.2 | 70.7 | 63.6 | 71.9 |
| + Self-consistent chain-of-thought | - | 73.2 | 75.7 | 81.7 | 69.7 | 64.1 | 72.9 |
| GPT-4 (gpt-4) | - | 75.0 | 43.0 | 73.0 | 57.0 | 77.0 | 65.0 |
| LLAMA2 (7B; Touvron et al., 2023) | - | 57.5 | 57.8 | 78.8 | 48.3 | 69.2 | 62.3 |
| LLAMA2 (13B; Touvron et al., 2023) | - | 55.9 | 67.3 | 80.2 | 50.3 | 72.8 | 65.3 |
| Mistral-v0.1 (7B; Jiang et al., 2023) | - | 51.0 | 59.6 | 83.0 | 42.9 | 75.3 | 62.4 |
| VERA-T5-xxl Liu et al. (2023) | ATOMIC | 71.2 | 61.7 | 76.4 | 57.7 | 67.5 | 66.9 |
| VERA-T5-xxl Liu et al. (2023) | ATM10X | 70.3 | 59.5 | 75.1 | 58.2 | 67.2 | 66.1 |
| VERA-T5-xxl Liu et al. (2023) | AbsATM | 73.2 | 63.0 | 77.2 | 58.1 | 68.1 | 68.0 |
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8 | 64.7 | 77.6 | 59.4 | 71.3 | 69.4 |
5.2.3 Zero-shot Commonsense QA
Task Setup.
The task of zero-shot commonsense QA involves selecting the most plausible option for commonsense questions without any supervision signals from benchmark data. Several methods have been proposed to tackle this task, including those by Shwartz et al. (2020); Bosselut et al. (2021); Kim et al. (2022); Shi et al. (2023). The most effective pipeline, as suggested by Ma et al. (2021), injects commonsense knowledge into language models via fine-tuning on QA pairs synthesized from knowledge in CSKBs. During the fine-tuning process, the head and relation of a triple are transformed into a question using natural language prompts, with the tail serving as the correct answer option. Distractors or negative examples are generated by randomly sampling tails from triples that do not share common keywords with the head. This fine-tuning procedure enhances the model’s knowledge not only for QA benchmarks constructed from CSKBs but also improves its ability to answer unseen commonsense questions in a more generalized manner.
We follow a similar approach to obtain our distilled models by first synthesizing QA pairs from various CSKBS. In addition to directly synthesizing from knowledge triples in ATOMIC, we augment ATOMIC by sampling triples from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations. The number of sampled triples is the same as in the original ATOMIC dataset. We then synthesize them into QA pairs, following the pipeline introduced by Ma et al. (2021). The resulting QA pairs are used to train different baseline models and CANDLE distilled models. For our distilled models, we utilize QA pairs sourced from CANDLE-instantiation augmented ATOMIC to train a DeBERTa-v3-large model using the marginal ranking loss Ma et al. (2021) and a T5-xxl model Raffel et al. (2020) following the training regime of VERA Liu et al. (2023). Following Ma et al. (2021), we evaluate the performance of the models on the validation split of five commonsense QA datasets: Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). Accuracy is used as the evaluation metric.
Baselines.
We compare our distilled models against various baselines. For pre-trained language models, we report the performances of vanilla RoBERTa-Large, DeBERTa-v3-Large, Self-talk Shwartz et al. (2020), COMET-DynaGen Bosselut et al. (2021), SMLM Banerjee and Baral (2020), MICO Su et al. (2022), MR Ma et al. (2021), STL-Adapter Kim et al. (2022), and the previous state-of-the-art CAR Wang et al. (2023a). For MR Ma et al. (2021) and CAR Wang et al. (2023a), DeBERTa-v3-Large is used as the backbone and their performances on ATOMIC-10X and AbstractATOMIC are also reported. For LLMs, we first report the performances of prompting GPT3.5 Brown et al. (2020), ChatGPT, GPT4 OpenAI (2023), LLAMA2, and Mistral in a zero-shot manner. For ChatGPT, advanced zero-shot prompting techniques such as chain-of-thought Wei et al. (2022) and self-consistency chain-of-thought Wang et al. (2023c) are used to assist the model in reasoning. We also train several VERA-T5-xxl baselines on different sets of QA pairs as LLM baselines.
Results and Analysis.
Table 4 presents the results, demonstrating that CANDLE distilled models generalize better than the baselines across several commonsense QA benchmarks. For instance, VERA demonstrates an average improvement of 1.4% compared to the best baseline. This can be attributed to the inclusion of new entities and events in CANDLE instantiations that are absent in other CSKBs, where CANDLE instantiations can aid in answering commonsense questions that require knowledge of these new instances. Furthermore, the distilled DeBERTa-v3-large model outperforms all baselines, including methods utilizing LLMs. By incorporating CANDLE instantiations, the performance of QA models is significantly enhanced, surpassing those utilizing ATOMIC-10X and AbstractATOMIC. This also shows the effectiveness of distilling knowledge with CANDLE in downstream tasks compared to simply applying symbolic knowledge distillation or conceptualization augmentation.
6 Conclusion
This paper introduces CANDLE, a distillation framework that realizes the chain of conceptualization and instantiation over CSKBs. We demonstrate the efficacy of CANDLE through comprehensive evaluations of the distilled knowledge and its positive impact on downstream tasks. Our research sheds light on leveraging LLMs to enable more robust and generalizable commonsense reasoning.
Limitations
The major limitation of CANDLE lies in the significant cost of distilling LLMs to obtain substantial knowledge. While the instantiation step of CANDLE utilizes the open foundation model LLAMA2, the conceptualization is still performed by ChatGPT due to the unsatisfactory performance of other LLMs and the superior quality of ChatGPT’s generation. Consequently, a considerable amount of funding is required to distill conceptualizations for CANDLE to function effectively. Future works should address this issue by exploring advanced LLMs or employing stronger open LLMs as the foundation for distilling conceptualizations, potentially utilizing advanced prompting techniques.
Furthermore, it should be noted that CANDLE has only been validated on the ATOMIC dataset. However, CANDLE is not limited to any specific format of commonsense knowledge, allowing it to operate on any CSKB. Future research can address this by extending the evaluation of CANDLE on other CSKBs Nguyen et al. (2023); Speer et al. (2017); Fang et al. (2023), and conducting follow-up experiments to explore their benefits on other downstream tasks or domains.
Ethics Statement
One of the main ethical concerns in this process is the potential for generating harmful or unethical content from models like ChatGPT and LLAMA2. To address this concern, we recruit four expert annotators, who are graduate and undergraduate students specializing in machine commonsense in natural language processing, to verify the ethics and potential harm of the generated content. A thorough assessment of a random sample has been conducted, and no significant harm has been identified. All training and evaluation datasets used are publicly available and shared under open-access licenses solely for research purposes, aligning with their intended usage. These datasets have been carefully anonymized and desensitized to protect data privacy and confidentiality. The expert annotators involved in this study are fully aware of the annotation protocol and the intended use of their annotations. Their participation in this research is voluntary, and they have agreed to contribute without receiving any compensation. Thus, the authors believe that this paper does not raise any ethical concerns to the best of their knowledge.
Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendices
Appendix A Distillation Details
This section provides additional details about the CANDLE distillation process not covered in the main body text. First, we present the prompts used to instruct ChatGPT to perform contextualized conceptualizations and LLAMA2 to perform contextualized instantiation. For prompting ChatGPT to distill conceptualizations, we use a few-shot prompt as shown below:
Following the given examples, you are required to conceptualize the instance (enclosed by []) in the last given event into abstract concepts. The concept should still fit into the instance’s original sentence. Make sure that the generated abstract concepts are general and not simply hypernyms of the instance.
Event <i>: PersonX enjoys drinking in the [bar], as a result, PersonX feels relaxed. [bar] can be conceptualized as Social Gathering Place
Event <N>: PersonX likes [painting on the beach], as a result, PersonX will go to the beach. [painting on the beach] can be conceptualized as
Similarly, for prompting LLAMA2-13B to distill instantiations based on previously generated conceptualizations, we use a few-shot prompt as shown below:
Following the given examples, you are required to instantiate the concept (enclosed by []) in the last given event into entities or events. If the event only contains the concept, then instantiate it to an event starting with a subject PersonX or PersonY. If the event contains other words, then instantiate it to an entity. The instance should still fit into the original sentence. Make sure that the generated instance is specific.
Event <i>: PersonX enjoys drinking in the [Social Gathering Place], as a result, PersonX feels relaxed. [Social Gathering Place] can be instantiated as beer festival
Event <N>: PersonX likes [exercise], as a result, PersonX will go to the stadium. [exercise] can be conceptualized as
| Abs.ATM | CANDLE | |
| #Unq. event | 15,388 | 15,359 |
| #Unq. instance | 21,493 | 21,442 |
| #Unq. conceptualization | 31,227 | 853,499 |
| #Tot. conceptualization | 503,588 | 6,181,391 |
| #Unq. instantiation | - | 676,737 |
| #Tot. instantiation | - | 6,181,391 |
| Avg. #concept/event | 32.73 | 173.33 |
| Avg. #Unq. concept/event | 28.33 | 167.76 |
| Avg. #concept/instance | 23.43 | 124.16 |
| Avg. #Unq. concept/instance | 17.27 | 100.88 |
| Relation | ATOMIC | Abs.ATM | CANDLE |
| xEffect | 78,832 | 938,330 | 964,765 |
| oEffect | 28,351 | 333,845 | 346,363 |
| xWant | 101,249 | 1,170,835 | 1,322,810 |
| oWant | 43,079 | 484,570 | 551,391 |
| xReact | 62,969 | 510,476 | 480,259 |
| oReact | 26,570 | 224,706 | 208,538 |
| xNeed | 74,272 | 900,429 | 894,338 |
| xAttr | 110,791 | 838,191 | 810,958 |
| xIntent | 45,490 | 519,813 | 601,969 |
| Total | 572,053 | 5,921,195 | 6,181,391 |
These prompts are consistent with our descriptions in Section 4.1 and Section 4.2, where the task description is first presented, followed by human-authored examples, and finally, the event we want to conceptualize or instantiate.
Additionally, we introduce some generation settings when prompting LLMs. For ChatGPT, we access it through the official OpenAI APIs111https://chat.openai.com/. The code of the accessed version is gpt-3.5-turbo-0613. We set the temperature to 0.7 and the maximum length for generated tokens to 200. To conceptualize all events in ATOMIC into 20 conceptualizations each, the time required for the distillation process is approximately ten days.
For LLAMA2, we access it via the Huggingface222https://huggingface.co/meta-llama/Llama-2-13b-chat-hf Library. The code of the accessed model is meta-llama/Llama-2-13b-chat-hf. When prompting, we use the Top-k sampling decoding strategy and set . We set the maximum length of generated tokens to 200. The models are hosted on sixteen NVIDIA-V100 GPUs, and the time required to distill the entire dataset is approximately one month.
After collecting 20 conceptualizations for every head event in ATOMIC and further instantiating them to new entities and events, we construct an expanded knowledge base of ATOMIC. We also include more statistics, as shown in Table 5 and Table 6. For instantiations, they share the same relational distribution as abstract commonsense triples since we only instantiate them once. These statistics indicate that, compared to AbstractATOMIC, which is the only available conceptualization benchmark based on ATOMIC, CANDLE contains more abstract commonsense triples and many more unique conceptualizations. According to our results, it can also be expected that the abstract knowledge distilled from CANDLE is of better quality than AbstractATOMIC, which human annotations or any filtering have not verified.
For critic filtering, we use the state-of-the-art conceptualization discriminator developed by Wang et al. (2023b). This discriminator is utilized to assess the plausibility of CANDLE distilled conceptualizations. It considers the original event, the instance being conceptualized, and the target concept as its inputs and generates a score ranging from 0 to 1 to represent plausibility. For instantiation, we use the pre-trained VERA model released by Liu et al. (2023). We convert the instantiated commonsense knowledge triple into a declarative statement and request an estimation of its plausibility from VERA. This estimation is provided as a score ranging from 0 to 1. The output scores from both models serve as the critical values assigned to each CANDLE distillation. These critical values are then subjected to further filtering based on various thresholds.
Furthermore, we employ BLEU soft uniqueness Zhu et al. (2018); West et al. (2022) to evaluate the distinctiveness of the newly introduced head events resulting from our chain of conceptualization and instantiation. To determine uniqueness, we define an instantiated head event, referred to as , as unique if , where represents the original head event in ATOMIC. The threshold of 0.5 is an empirical threshold. Our empirical results demonstrate that 78.6% of the instantiated events are unique compared to ATOMIC, highlighting the effectiveness of CANDLE in enhancing the semantic coverage of the CSKB.
Future works can experiment with applying CANDLE distillation framework to other downstream tasks, such as complex-query answering Bai et al. (2023), script generation Yuan et al. (2023), pragmatics Ou et al. (2023a, b), analogical reasoning Cheng et al. (2023), and E-Commerce understanding Yu et al. (2023).
Appendix B Dataset Descriptions
This section covers additional details and statistics of datasets and benchmarks used in downstream task evaluations. Please refer to the corresponding subsection for each downstream task.
| Data | Type | Train | Dev | Test |
| #event | 107,384 | 12,117 | 11,503 | |
| #triple | 65,386 | 8,403 | 7,408 | |
| #event | 304,983 | 36,023 | 31,578 | |
| #triple | 4,851,272 | 499,523 | 570,400 |
B.1 CSKB Conceptualization
In CSKB Conceptualization tasks, we use the AbstractATOMIC He et al. (2022a) dataset as the evaluation benchmark. It is a benchmark dataset built upon ATOMIC and consists of event conceptualization data and abstract knowledge triples. The event conceptualizations are based on head events in ATOMIC, identified through syntactic parsing and matching with rules to search for concept candidates in Probase Wu et al. (2012) and WordNet Miller (1995). The abstract knowledge triples connect conceptualized head events with their non-abstract counterparts from ATOMIC, forming commonsense knowledge at the concept level. Human annotations are used to verify the correctness of some conceptualizations and their resulting abstract commonsense triples. In total, 131K conceptualizations of 7K (45%) ATOMIC head events and 81K (1.3%) conceptualized triples are manually annotated, with a large number remaining unlabeled. Detailed statistics are shown in Table 7. In this paper, we evaluate all models using the test set from the annotated subset as the evaluation data. Meanwhile, we obtain CANDLE distilled models using the training set from the annotated subset to fine-tune discriminative models pre-trained on CANDLE conceptualizations.
B.2 Generative Commonsense Inference
To evaluate COMET, we adopt the same evaluation setting employed by Hwang et al. (2021) for assessing commonsense generative models on the ATOMIC dataset’s test set. We use the entire test set, consisting of 34,689 triples across 23 different commonsense relations, to ensure the robustness of the evaluation. Additionally, we use the full training set to fine-tune models that were pretrained on various CSKBs and CANDLE instantiations.
Recently, West et al. (2023) successfully trained a powerful commonsense inference generator using an open-format symbolic knowledge distillation framework. Once they release their data and models, we will incorporate them as another baseline in our comparisons.
| aNLI | CSQA | PIQA | SIQA | WG | |
| #QA Pairs | 1,532 | 1,221 | 1,838 | 1,954 | 1,267 |
| #Options | 2 | 5 | 2 | 3 | 2 |
B.3 Zero-shot Commonsense QA
We follow Ma et al. (2021); Wang et al. (2023a); Shi et al. (2023) and use the validation split of five commonsense QA benchmarks: Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). These benchmarks evaluate different aspects, including abductive reasoning, concept-level commonsense reasoning, physical commonsense understanding, emotional and social commonsense reasoning, and pronoun resolution. The validation splits are used as the official test sets may not be publicly available. Statistics on the number of QA pairs and the number of options per question are reported in Table 8.
| Task | Prompt | |||||||
| Event. |
| Triple. |
| COMET |
| aNLI |
| CSQA |
| PIQA |
| SIQA |
| WG |
Appendix C Implementation Details
This section provides additional information about implementation details in downstream task evaluations.
First, we use the Huggingface Library Wolf et al. (2020) to build all models. We reproduce all baselines according to implementation details described in their original papers. The reported results are consistent with their original papers if the same experiment is included. For CANDLE distilled models, please refer to the subsections (Appendix C.1, C.2, C.3) below.
For methods involving LLMs, we use their instruction fine-tuned versions as the backbone for the baselines. For LLAMA2, the accessed version is meta-llama/Llama-2-7b/13b-chat-hf. For Mistral, we use mistralai/Mistral -7B-Instruct-v0.1. This remains consistent whether we prompt them directly or fine-tune them for downstream tasks, as we have observed that the instruction-finetuned versions generally result in better performance. For ChatGPT, we access it through Microsoft Azure APIs333https://azure.microsoft.com/en-us/products/ai-services/. The code of the accessed version for ChatGPT is gpt-35-turbo-20230515, and for GPT4 is gpt-4-20230515. The maximum generation length is set to 100 tokens for all tasks. For fine-tuning LLAMA2 and Mistral, we use the open code base of LLaMa-Factory444https://github.com/hiyouga/LLaMA-Factory. Please refer to the subsections below for hyperparameter settings.
All experiments are conducted on sixteen NVIDIA-V100 (32G) GPUs.
For baselines involving prompting LLMs, we follow the approach done by Robinson and Wingate (2023) and Chan et al. (2023), where each task is formulated in either a generative format or as multiple-choice QA. Table 9 shows the prompts used in zero-shot prompting scenarios. To incorporate five-shot exemplars, we include five randomly selected examples from the training set of each benchmark. These examples are merged into the prompt using the same format as the question, with the addition of including the answer at the end. For chain-of-thought reasoning, we prompt LLM in a two-step inference process. In the first step, we delve deeper into the question by requesting an intermediate-step rationale. Then, in the second step, we seek an answer based on the question and the previous step’s response by asking LLM to answer “Yes or No only” or select the correct option from a set of answers directly.
| Training Data | Bleu-1 | Bleu-2 | Bleu-3 | Bleu-4 | METEOR | ROUGE-L | CIDEr | BERTScore |
| Backbone: GPT2-XL Radford et al. (2019) 1.5B | ||||||||
| Zero-shot | 4.350 | 1.598 | 0.732 | 0.293 | 5.702 | 5.030 | 0.792 | 37.11 |
| ATOMIC | 32.23 | 19.06 | 13.27 | 10.28 | 17.63 | 25.50 | 20.15 | 58.39 |
| + Finetune | 45.72 | 29.18 | 21.12 | 16.15 | 29.97 | 49.69 | 64.61 | 76.09 |
| ATOMIC | 42.15 | 25.77 | 17.82 | 13.14 | 29.82 | 47.61 | 63.70 | 70.39 |
| ATOMIC-10X | 33.69 | 18.82 | 11.71 | 7.910 | 18.78 | 25.69 | 19.29 | 61.47 |
| + Finetune | 45.38 | 29.20 | 21.09 | 16.15 | 30.09 | 49.86 | 65.02 | 75.89 |
| AbstractATOMIC | 29.46 | 17.16 | 11.89 | 9.019 | 17.42 | 24.30 | 19.95 | 57.83 |
| + Finetune | 45.30 | 29.08 | 21.00 | 16.06 | 29.98 | 48.61 | 63.98 | 75.56 |
| CANDLE Distilled | 26.91 | 16.44 | 12.31 | 10.28 | 17.66 | 23.66 | 21.36 | 57.15 |
| + Finetune | 50.71 | 33.85 | 25.55 | 20.43 | 32.45 | 51.91 | 69.68 | 76.86 |
| Backbone: ChatGPT OpenAI (2022) (openai/gpt-3.5-turbo) | ||||||||
| Zero-shot | 11.82 | 4.258 | 1.891 | 0.926 | 13.87 | 13.73 | 4.350 | 49.28 |
| Five-shot | 26.32 | 12.50 | 7.160 | 4.415 | 18.60 | 24.65 | 8.313 | 58.69 |
| Chain-of-thought | 9.906 | 3.568 | 1.556 | 0.736 | 11.85 | 11.02 | 2.905 | 46.17 |
| Backbone: LLAMA2 Touvron et al. (2023) 7B | ||||||||
| Zero-shot | 18.26 | 7.453 | 3.594 | 1.945 | 15.90 | 20.28 | 8.872 | 48.23 |
| Five-shot | 31.22 | 16.87 | 9.767 | 5.989 | 19.74 | 27.67 | 17.83 | 58.41 |
| ATOMIC | 29.94 | 16.44 | 10.03 | 6.631 | 19.02 | 25.75 | 18.71 | 59.68 |
| + Finetune | 42.04 | 23.01 | 14.10 | 9.125 | 27.80 | 42.90 | 53.17 | 71.52 |
| ATOMIC | 41.07 | 22.46 | 13.62 | 8.619 | 27.74 | 42.42 | 53.28 | 71.77 |
| ATOMIC-10X | 33.06 | 17.65 | 9.986 | 6.078 | 19.22 | 25.32 | 17.80 | 61.25 |
| + Finetune | 42.07 | 23.08 | 14.14 | 9.198 | 28.14 | 42.75 | 53.69 | 71.93 |
| AbstractATOMIC | 26.08 | 13.27 | 7.799 | 5.018 | 15.08 | 21.20 | 14.78 | 56.83 |
| + Finetune | 42.78 | 23.64 | 14.58 | 9.471 | 27.74 | 42.55 | 53.12 | 71.51 |
| CANDLE Distilled | 28.93 | 15.56 | 9.468 | 6.140 | 18.60 | 25.37 | 17.20 | 60.27 |
| + Finetune | 43.86 | 24.40 | 15.12 | 10.00 | 28.36 | 43.86 | 54.25 | 72.94 |
C.1 CSKB Conceptualization
For RoBERTa and DeBERTa-v3, we use a learning rate of 5e-6 and a batch size of 64. To optimize the models, we use an AdamW optimizer Loshchilov and Hutter (2019) and evaluate the model’s performance every 25 steps. The maximum sequence lengths for the tokenizers are set to 25 and 35 for the two discriminative subtasks, respectively. Early stopping is used where the best checkpoint is selected when the largest validation accuracy is achieved. The models are trained on CANDLE distillation for one epoch and fine-tuned on the training set of AbstractATOMIC for one epoch.
For LLMs, such as LLAMA2 and Mistral, we use LoRA for fine-tuning, and the LoRA rank and are set to 64 and 64. We use an Adam Kingma and Ba (2015) optimizer with a learning rate of 5e-6 and a batch size of 64. The models are fine-tuned for two epochs, and the checkpoint with the highest validation set accuracy is selected. All experiments are repeated three times using different random seeds, and the average performances are reported.
For VERA, we follow the exact same implementation555https://github.com/liujch1998/vera as released by Liu et al. (2023). To transform our binary classification subtasks into declarative formats, we begin by converting each piece of data into a declarative sentence using predefined natural language templates. Next, we create a corresponding negative statement by simply incorporating the word “not” into the correct sentence. For instance, a pair of statements is: “PersonX enjoys drinking at the bar. Bar is a social gathering place.” and “PersonX enjoys drinking at the bar. Bar is not a social gathering place.” The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation. The model is trained on CANDLE distillation for one epoch and then fine-tuned on AbstractATOMIC for another.
C.2 Generative Commonsense Inference
For COMET, we implement the open-sourced code by Hwang et al. (2021) as our base to fine-tune the GPT2 model. The model is first pre-trained on CANDLE instantiations for one epoch, followed by fine-tuning on ATOMIC for another epoch. An Adam Kingma and Ba (2015) optimizer is used with a learning rate of 1e-5 and a batch size of 32. A linear scheduler is used to decrease the learning rate gradually.
For LLAMA2-7B, we fine-tune it with the DeepSpeed framework Aminabadi et al. (2022) by using FP16 as the precision. We optimize the model with an Adam Kingma and Ba (2015) optimizer with a learning rate of 1e-4 and a batch size of 64. The maximum length for the input and generated sentence concatenation is 500. We warm up the model with 3000 steps and evaluate the model every 1000 steps. A linear scheduler is also used. The LoRA rank is set to 8, and the is set to 32.
In Table 10, we present supplementary automatic evaluation results, including models that have been pre-trained solely on CSKBs and CANDLE instantiations without subsequent fine-tuning.
| Original | Concept./Instant. | Critic |
| PersonX swims in the lake, as a result, PersonX feels, tired. | PersonX swims in freshwater, as a result, PersonX feels, tired. | 0.97 |
| PersonX swims in the sea, as a result, PersonX feels, tired. | 0.87 | |
| PersonX swims, as a result, PersonX feels, tired. | 0.89 | |
| PersonX swims every week, as a result, PersonX feels, tired. | 0.81 | |
| PersonX is sitting in class, as a result, PersonX will, learns something. | PersonX is sitting in instructional period, as a result, PersonX will, learns something. | 0.54 |
| PersonX is sitting in a math class, as a result, PersonX will, learns something. | 0.75 | |
| PersonX study, as a result, PersonX will, learns something. | 0.78 | |
| PersonX learns how to do the exam, as a result, PersonX will, learns something. | 0.81 | |
| PersonX buys PersonY a gift, as a result, PersonY feels, joyful. | remembrance, as a result, PersonY feels, joyful. | 0.19 |
| PersonX reminisce, as a result, PersonY feels, joyful. | 0.27 | |
| PersonX shopping, as a result, PersonY feels, joyful. | 0.61 | |
| PersonX buys a new toy for PersonY, as a result, PersonY feels, joyful. | 0.90 | |
| PersonX always fought, as a result, PersonY feels, angry. | PersonX always violent behavior, as a result, PersonY feels, angry. | 0.98 |
| PersonX always punch others hardly, as a result, PersonY feels, angry. | 0.91 | |
| combative personality, as a result, PersonY feels, angry. | 0.98 | |
| PersonX PersonX likes to join a fight, as a result, PersonY feels, angry. | 0.85 | |
| PersonX gets a new bike, as a result, PersonX wants, to ride it. | PersonX gets a transportation tool, as a result, PersonX wants, to ride it. | 0.92 |
| PersonX gets a motor, as a result, PersonX wants, to ride it. | 0.98 | |
| bike possession, as a result, PersonX wants, to ride it. | 0.93 | |
| PersonX has a nice bicycle, as a result, PersonX wants, to ride it. | 0.89 | |
| PersonX spends time with PersonY, PersonX is seen as, social. | PersonX spends love-building period with PersonY, PersonX is seen as, social. | 0.05 |
| PersonX spends time in love with PersonY, PersonX is seen as, social. | 0.37 | |
| social activity, PersonX is seen as, social. | 0.64 | |
| PersonX enjoys going to parties, PersonX is seen as, social. | 0.73 | |
| PersonX hears sirens, as a result, PersonX will, make way to the siren. | emergency response, as a result, PersonX will, make way to the siren. | 0.37 |
| PersonX sees an ambulance coming, as a result, PersonX will, make way to the siren. | 0.74 | |
| PersonX hears loud noise, as a result, PersonX will, make way to the siren. | 0.67 | |
| PersonX hears a fire truck beeping, as a result, PersonX will, make way to the siren. | 0.77 |
| Model/Method | CSKB | a-NLI | CSQA | PIQA | SIQA | WG | Avg. |
| Pre-trained Language Models | |||||||
| Random Vote | - | 50.0 | 20.0 | 50.0 | 33.3 | 50.0 | 40.7 |
| Majority Vote | - | 50.8 | 20.9 | 50.5 | 33.6 | 50.4 | 41.2 |
| GPT2-L Radford et al. (2019) | - | 56.5 | 41.4 | 68.9 | 44.6 | 53.2 | 52.9 |
| RoBERTa-L Liu et al. (2019) | - | 65.5 | 45.0 | 67.6 | 47.3 | 57.5 | 56.6 |
| DeBERTa-v3-L He et al. (2023) | - | 59.9 | 25.4 | 44.8 | 47.8 | 50.3 | 45.6 |
| Self-talk Shwartz et al. (2020) | - | - | 32.4 | 70.2 | 46.2 | 54.7 | - |
| SMLM Banerjee and Baral (2020) | * | 65.3 | 38.8 | - | 48.5 | - | - |
| COMET-DynGen Bosselut et al. (2021) | ATOMIC | - | - | - | 50.1 | - | - |
| MICO Su et al. (2022) | ATOMIC | - | 44.2 | - | 56.0 | - | - |
| STL-PLM Kim et al. (2022) | ATOMIC | 71.6 | 64.0 | 72.2 | 63.2 | 60.5 | 66.3 |
| MTL Kim et al. (2022) | CWWV | 69.6 | 67.3 | 72.5 | 52.0 | 57.2 | 63.7 |
| MTL Kim et al. (2022) | CSKG | 69.8 | 67.1 | 72.0 | 61.9 | 59.3 | 66.0 |
| STL-Adapter Kim et al. (2022) | ATOMIC | 71.3 | 66.5 | 71.1 | 64.4 | 60.3 | 66.7 |
| STL-Adapter Kim et al. (2022) | CSKG | 71.5 | 66.7 | 72.1 | 64.7 | 59.0 | 66.8 |
| RoBERTa-L (MR) Ma et al. (2021) | ATM | 70.8 | 64.2 | 71.7 | 61.0 | 60.7 | 65.7 |
| RoBERTa-L (MR) Ma et al. (2021) | ATOMIC | 70.8 | 64.2 | 72.1 | 63.1 | 59.2 | 65.9 |
| RoBERTa-L (MR) Ma et al. (2021) | CWWV | 70.0 | 67.9 | 72.0 | 54.8 | 59.4 | 64.8 |
| RoBERTa-L (MR) Ma et al. (2021) | CSKG | 70.5 | 67.4 | 72.4 | 63.2 | 60.9 | 66.8 |
| DeBERTa-v3-L (MR) Ma et al. (2021) | ATM10X | 75.1 | 71.6 | 79.0 | 59.7 | 71.7 | 71.4 |
| DeBERTa-v3-L (MR) Ma et al. (2021) | ATOMIC | 76.0 | 67.0 | 78.0 | 62.1 | 76.0 | 71.8 |
| ZS-Fusion Kim et al. (2022) | CWWV | 69.6 | 67.6 | 73.1 | 53.7 | 59.5 | 64.7 |
| ZS-Fusion Kim et al. (2022) | CSKG | 72.4 | 68.3 | 73.0 | 66.7 | 60.9 | 68.3 |
| MKIF Guan et al. (2023) | CSKG | 72.5 | 71.0 | 73.1 | - | 61.0 | - |
| CAR-RoBERTa-L Wang et al. (2023a) | ATOMIC | 72.3 | 64.8 | 73.2 | 64.8 | 61.3 | 67.3 |
| CAR-RoBERTa-L Wang et al. (2023a) | AbsATM | 72.7 | 66.3 | 73.2 | 64.0 | 62.0 | 67.6 |
| CAR-DeBERTa-v3-L Wang et al. (2023a) | ATOMIC | 78.9 | 67.2 | 78.6 | 63.8 | 78.1 | 73.3 |
| CAR-DeBERTa-v3-L Wang et al. (2023a) | AbsATM | 79.6 | 69.3 | 78.6 | 64.0 | 78.2 | 73.9 |
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2 | 69.9 | 80.3 | 65.9 | 78.3 | 74.9 |
| Large Language Models | |||||||
| GPT-3.5 (text-davinci-003) | - | 61.8 | 68.9 | 67.8 | 68.0 | 60.7 | 65.4 |
| ChatGPT (gpt-3.5-turbo) | - | 69.3 | 74.5 | 75.1 | 69.5 | 62.8 | 70.2 |
| + Chain-of-thought | - | 70.5 | 75.5 | 79.2 | 70.7 | 63.6 | 71.9 |
| + Self-consistent chain-of-thought | - | 73.2 | 75.7 | 81.7 | 69.7 | 64.1 | 72.9 |
| GPT-4 (gpt-4) | - | 75.0 | 43.0 | 73.0 | 57.0 | 77.0 | 65.0 |
| LLAMA2 (7B; Touvron et al., 2023) | - | 57.5 | 57.8 | 78.8 | 48.3 | 69.2 | 62.3 |
| LLAMA2 (13B; Touvron et al., 2023) | - | 55.9 | 67.3 | 80.2 | 50.3 | 72.8 | 65.3 |
| Mistral-v0.1 (7B; Jiang et al., 2023) | - | 51.0 | 59.6 | 83.0 | 42.9 | 75.3 | 62.4 |
| VERA-T5-xxl Liu et al. (2023) | ATOMIC | 71.2 | 61.7 | 76.4 | 57.7 | 67.5 | 66.9 |
| VERA-T5-xxl Liu et al. (2023) | ATM10X | 70.3 | 59.5 | 75.1 | 58.2 | 67.2 | 66.1 |
| VERA-T5-xxl Liu et al. (2023) | AbsATM | 73.2 | 63.0 | 77.2 | 58.1 | 68.1 | 68.0 |
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8 | 64.7 | 77.6 | 59.4 | 71.3 | 69.4 |
| Supervised Learning & Human Performance | |||||||
| RoBERTa-L (Supervised) | - | 85.6 | 78.5 | 79.2 | 76.6 | 79.3 | 79.8 |
| DeBERTa-v3-L (Supervised) | - | 89.0 | 82.1 | 84.5 | 80.1 | 84.1 | 84.0 |
| VERA-T5 (Multitask Supervised) | - | 83.9 | 77.8 | 88.5 | 80.1 | 92.4 | 84.5 |
| Human Performance | - | 91.4 | 88.9 | 94.9 | 86.9 | 94.1 | 91.2 |
C.3 Zero-shot Commonsense QA
For the task of zero-shot commonsense QA, we adopt the code base provided by Wang et al. (2023a)666https://github.com/HKUST-KnowComp/CAR and Liu et al. (2023)777https://github.com/liujch1998/vera to train two CANDLE distilled models. All hyperparameters are kept unchanged from their original implementations as default settings. The models are trained for two epochs using QA pairs obtained from augmented-ATOMIC, including augmentations from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations.
Meanwhile, we present a comprehensive table presenting the results of all current methodologies for the task of zero-shot commonsense QA in Table 12. Notably, our CANDLE distilled models continue to exhibit strong performance compared to other models pre-trained on QA pairs sourced from multiple CSKBs. This serves as compelling evidence for the efficacy of CANDLE.
Appendix D Annotation Details
This paper utilizes expert annotations to assess the quality of distilled conceptualizations and instantiations, as well as evaluate the generations of different models for the COMET downstream task. Four graduate students with ample experience in natural language processing research and expertise in commonsense reasoning are recruited as expert annotators to carry out the annotations. Their participation in the annotation process is voluntary and unpaid, in accordance with local laws, and is considered a contribution to this paper. Detailed instructions are provided to the annotators for each task, ensuring that they understand the requirements thoroughly. For each question, we also highlight the part to be considered by the annotators for their convenience. The annotation process is conducted independently, without any internal discussions among the annotators regarding the results. For each task, two annotators independently vote for each triple, and only when both annotators provide a positive vote will the triple be considered accepted or plausible. To prevent bias and ensure impartial results for CANDLE, the task input is randomly shuffled during the annotation process. As a result, the expert annotators achieve a pairwise agreement (IAA; Landis and Koch, 1977) of 0.80 and a Fleiss-kappa Fleiss (1971) of 0.61, indicating a remarkably high level of internal agreement.
Appendix E Case Study
We present some examples in Table 11 to show conceptualizations and instantiations generated by CANDLE, along with their corresponding critic values assigned by our filtering discriminators. Most of the generations are of high quality, as expected, and can introduce novel events compared to the original knowledge.